### v2

@@ 1 Introduction
Modern large-scale language models (Brown et al. 2020, Chowdhery et al. 2023, OpenAI 2023, Kaplan et al. 2020, Wei et al. 2022) have revolutionized artificial intelligence by demonstrating unprecedented capabilities in natural language generation across diverse linguistic domains and situational contexts. These sophisticated neural networks, trained on extensive corpora of textual data, now serve as foundational components for numerous real-world applications. Their deployment spans conversational agents (Anthropic 2023, Character 2021, OpenAI 2019, 2023), information retrieval systems (Microsoft 2023, Google 2023, Komo 2023, Perplexity 2022, You.com 2020), programming aids (Amazon 2023, GitHub 2021, Replit 2018), and even clinical decision support tools (Cascella et al. 2023, Peng et al. 2023, Sallam 2023), demonstrating remarkable versatility in professional and consumer domains alike.
@@ 1.1 Main Contribution
Negative Results for Existing Scheduling Algorithms and NP-hardness. In Section 2, we first establish fundamental limitations of current methods by analyzing the model in Jaillet et al. (2025) under realistic conditions. Our key theoretical findings reveal that: (1) without uniform input sizes, both first-come-first-serve and shortest-first scheduling yield unbounded competitive ratios (regardless of using output length or total sequence length as the metric), and (2) the general scheduling problem is NP-hard. These results formally justify the challenge need for new scheduling approaches.
@@
Approximation Algorithms for Practical Deployment. Since scheduling latency directly impacts overall inference time in production systems, we investigate efficient approximations of Sorted-F that maintain its performance benefits while reducing computational overhead. The primary complexity bottleneck in Sorted-F lies in its batch partitioning phase based on our quality metric. In Section 5, we analyze four approximation approaches: (1) exact dynamic programming, (2) scaled dynamic programming, (3) local swap search, and (4) quantile greedy selection. For each method, we characterize both its theoretical computational complexity and practical applicability, specifying the problem scales where each variant proves to be most effective. These analyses provide system designers with clear guidelines for algorithm selection based on their specific latency requirements and workload characteristics.
@@
Numerical Experiments. In Section 7, we present comprehensive empirical evaluations using real-world data to validate our theoretical findings. To accurately capture the practical challenges of variable-length inputs, we construct a mixed dataset combining: (1) short conversational prompts from Zheng et al. (2023), and (2) long document summarization tasks from Cohan et al. (2018). This experimental design reflects realistic deployment scenarios where LLM servers must handle diverse request types simultaneously. We evaluate four scheduling approaches: first-come-first-serve, shortest-first, our LP-based Sorted-LP, and our main proposed algorithm Sorted-F. The results demonstrate that Sorted-F achieves significantly lower average latency, confirming both the theoretical robustness and practical effectiveness of our approach.
@@ 1.2 Other Related Works
In contrast, a few recent studies aim to develop theoretical foundations for LLM inference. The primary mathematical model considered in this paper builds on the framework introduced by Jaillet et al. (2025). Ao et al. (2025) examine inference scheduling under multiple objectives and propose algorithms with provable performance guarantees. Additionally, Li et al. (2025) analyze the stability conditions of LLM inference when operating on a single computational worker.
@@ 2 Model Review
In this section, we review the model introduced by Jaillet et al. (2025). The setting considers LLM inference on a single computational worker with a key-value (KV) cache limit $M>0$, meaning that the worker can hold at most $M$ tokens in memory at any given time. In practice, $M$ depends on the size of the model and the hardware specifications (e.g., GPU memory), and is assumed to be known to the decision-maker.
We consider a discrete-time system where $n$ prompts arrive simultaneously at time $t=0$. Each prompt $i\in[n]$ is characterized by a pair $(s_{i},o_{i})$, where $s_{i}$ denotes the input size (i.e., the number of tokens in the input prompt), and $o_{i}$ denotes the number of output tokens to be generated. We assume that $o_{i}$ can be accurately predicted, and that both scale as $s_{i},o_{i}=\epsilon(M)$, where $\epsilon(M)$ is a term asymptotically smaller than $M$. This captures the natural regime where individual requests consume a vanishing fraction of total memory.
Each request $i$ must be processed token by token in order. The memory required to generate the $j$th token (where $j\in[o_{i}]$) is $s_{i}+j$, reflecting the growing size of the KV cache as new tokens are appended. At each time step $t$, the system processes a batch of tokens $\mathcal{I}_{t}$, which may consist of tokens from different prompts, but cannot include multiple tokens from the same prompt, as token generation is sequential. To ensure feasibility, the total memory usage in any batch must not exceed $M$, i.e.,
@@
where $a_{i}$ denotes the index of the token from request $i$ included in the batch.
Our objective is to minimize the total end-to-end latency, defined as the sum of completion times for all requests. Since all prompts arrive at $t=0$, the end-to-end latency of request $i$ is equal to the time at which its final token is processed. Letting $c_{i}(\Lambda)$ denote the completion time of request $i$ under schedule $\Lambda$, the total end-to-end latency is
@@
To compute the optimal schedule, Jaillet et al. (2025) proposes an integer linear program (ILP). However, the ILP is computationally intractable in real-time inference settings, where scheduling decisions must be made within milliseconds.
Thus, our goal is to design efficient approximation algorithms. Let Optimal denote the optimal schedule with the smallest end-to-end latency for request set $\mathcal{I}$. For any scheduling algorithm $\Lambda$, we evaluate its performance using the competitive ratio:
@@
The competitive ratio measures the worst-case performance of the algorithm relative to the optimum. It is always at least 1, with smaller values indicating better performance. Our objective is to develop algorithms with low competitive ratios that are practical for real-time deployment.
2.1 Scheduling Challenges with Long Prompts
In Jaillet et al. (2025), the authors propose a memory-constrained shortest-first scheduling algorithm (MC-SF). The algorithm prioritizes requests with smaller output lengths $o_{i}$. When considering whether to add a prompt to the current batch, the algorithm precomputes the memory usage at each future time step to determine whether the addition would violate the memory constraint at any point in time. Formally, suppose a subset $U\subset R_{t}$ is considered for inclusion in the batch at time $t$. Define $t_{\max}(U):=\max_{i\in U}\{t+o_{i}\}$ as the latest possible completion time for jobs in $U$, assuming processing begins at time $t$. To ensure that the memory constraint is satisfied throughout the interval $t,t_{\max}(U)]$, the algorithm verifies that the total memory usage at each intermediate time $t^{\prime}\in[t,t_{\max}(U)]$ remains within the KV cache capacity $M$. Specifically, the memory-constrained shortest-first algorithm checks the following condition to determine whether including the batch $U$ is feasible at time $t$:
$\sum_{i\in S^{(t)}}(s+t^{\prime}-p_{i})\cdot\mathbbm{1}_{{o_{i}\geq t^{\prime}-p_{i}}}+\sum_{i\in U}(s+t^{\prime}-t)\cdot\mathbbm{1}_{{o_{i}\geq t^{\prime}-t}}\leq M,\quad\forall t^{\prime}\in[t,t_{\max}(U)].$ (1)
Here, $S^{(t)}$ denotes the set of jobs already in progress at time $t$, and $p_{i}$ represents the most recent start time of job $i$. The first summation accounts for the memory consumption of ongoing jobs that may still be active at time $t^{\prime}$, while the second summation captures the memory usage of new jobs in $U$ starting at time $t$. The indicator functions ensure that only those tokens still being generated at time $t^{\prime}$ contribute to the total memory usage.
Jaillet et al. (2025) demonstrates the strong theoretical and empirical performance of Algorithm MC-SF under the simplifying assumption that $s_{i}=s$ for all $i\in[n]$. This assumption makes the Algorithm MC-SF particularly natural: since $s_{i}$ determines the base memory usage, $o_{i}$ governs the processing time, and $s_{i}+o_{i}$ defines the peak memory usage, requests with smaller $o_{i}$ require less peak memory and complete faster. Thus, under uniform $s_{i}$, prioritizing shorter jobs both improves throughput and better utilizes memory.
However, the assumption of constant $s_{i}$ may not hold in practice, especially when the system must handle a mix of long and short prompts, leading to significant variability in input sizes. Relaxing this assumption introduces substantial complexity. For instance, consider two requests $i$ and $j$ where $s_{i}<s_{j}$ but $o_{i}>o_{j}$: request $i$ consumes less memory per time unit but runs longer, while request $j$ uses more memory but finishes quickly. In such cases, it is unclear which request should be prioritized, as the trade-off between memory consumption and completion time becomes nontrivial. Next, we formally show that when $s_{i}$ is allowed to vary, the competitive ratio of Algorithm MC-SF becomes unbounded.
@@ 2.1 Scheduling Challenges with Long Prompts
Suppose that each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$ and $s_{i}+o_{i}\leq M$, then Algorithm MC-SF achieves an unbounded competitive ratio. Specifically, as $M\to\infty$, we have
$\mathrm{CR}(\textup{{MC-SF}})\to\infty.$
The proof of Theorem 2.1 can be found in Appendix 8. Theorem 2.1 illustrates that relaxing the assumption $s_{i}=s$ for all $i\in[n]$ can lead to many instances where existing scheduling algorithms perform poorly. In Appendix 8, we also show that if we do shortest-first according to $s_{i}+o_{i}$, the competitive ratio is still unbounded. Furthermore, the following theorem establishes that, without the assumption $s_{i}=s$, the problem of minimizing total end-to-end latency is NP-hard [^1] It remains an open question that either minimizing total end-to-end latency is NP-hard or not under the assumption $s_{i}=s$..
@@
Under the model suggested in Jaillet et al. (2025), suppose that each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$ and $s_{i}+o_{i}\leq M$, then minimizing the total end-to-end latency is NP-hard.
We prove Theorem 2.2 by reduce the problem to a 3-Partition problem, and the rigorous proof can be found in Appendix 8. For readers interested in alternative objectives, we include in Appendix 8 a result (Theorem 8.5) showing that, under the same model, the problem remains NP-hard when the objective is to minimize the makespan. While this result is not central to the main focus of the paper, it further highlights the computational challenges of scheduling in this setting.
@@ 3 Algorithm Description
where $\mathcal{X}$ is a set of requests, and $|\mathcal{X}|$ is the cardinality of the request set.
Smaller values of $F(\mathcal{X})$ indicate higher scheduling priority. This metric fundamentally improves upon shortest-first methods by naturally balancing batch size and response lengths. To illustrate its operational intuition, we begin with a concrete numerical example that demonstrates how Sorted-F outperforms shortest-first scheduling (MC-SF):
@@
Under MC-SF (which prioritizes shorter outputs), Type 1 would be scheduled first since $o_{1}=1<o_{2}=2$. After Type 1 completes at time $t=1$, Type 2 requests are processed in batches. With memory constraint $M=64$, each batch can accommodate $\lfloor 64/(1+2)\rfloor=21$ Type 2 requests simultaneously. The total end-to-end latency (TEL) is
@@
Building on this concrete demonstration, we now generalize the intuition through a symbolic example that shows why the F-metric naturally leads to optimal scheduling decisions:

### v3

@@
We study offline scheduling for large language model (LLM) serving under a fixed KV-cache memory budget, where requests have heterogeneous prompt (prefill) and response (decode) lengths. Prompt tokens determine initial KV usage, and each generated token increases memory by one unit. Given a backlog of n requests arriving together, we schedule mixed prefill and decode batches to minimize total end-to-end latency. We show that heterogeneity in prompt lengths makes the problem computationally intractable and that widely used heuristics such as first-come-first-served and shortest-first can be arbitrarily suboptimal. We propose Sorted-F, which repeatedly forms feasible batches using a new selection metric that balances batch size against downstream decode cost, and prove it achieves a constant-factor guarantee on total latency. We further develop practical variants—an exact solver for small instances and fast heuristics for larger ones—and evaluate them on a public workload spanning short conversations and long-document summarization, where they consistently reduce average latency relative to standard baselines. Our results highlight that during peak-hour tidal backlogs, greedy GPU packing or short-request prioritization can perform poorly when prompt lengths vary widely, and provide a principled, tunable framework for designing production batch schedulers and planning capacity in memory-constrained LLM serving systems.
@@ 1 Introduction
Modern large-scale language models (Brown et al. 2020, Chowdhery et al. 2023, OpenAI 2023) have revolutionized artificial intelligence by demonstrating unprecedented capabilities in natural language generation across diverse linguistic domains and situational contexts. These sophisticated neural networks, trained on extensive corpora of textual data, now serve as foundational components for numerous real-world applications. Their deployment spans conversational agents (Anthropic 2023, OpenAI 2023), information retrieval systems (Microsoft 2023, Google 2023), programming aids (Amazon 2023, GitHub 2021), and even clinical decision support tools (Cascella et al. 2023, Peng et al. 2023), demonstrating remarkable versatility in professional and consumer domains alike.
@@ 1.1 Main Contribution
Negative Results for Existing Scheduling Algorithms and NP-hardness. In Section 2, we first establish fundamental limitations of current methods by analyzing the model in Jaillet et al. (2025) under realistic conditions. Our key theoretical findings reveal that: (1) without uniform input sizes, both first-come-first-served and shortest-first scheduling yield unbounded competitive ratios (regardless of using output length or total sequence length as the metric), and (2) the general scheduling problem is NP-hard. These results formally justify the need for new scheduling approaches.
@@
Approximation Algorithms for Practical Deployment. Since scheduling latency directly impacts overall inference time in production systems, we investigate efficient approximations of Sorted-F that maintain its performance benefits while reducing computational overhead. The primary complexity bottleneck in Sorted-F lies in its batch partitioning phase based on our quality metric. In Section 5, we analyze three approximation approaches: (1) exact dynamic programming, (2) local swap search, and (3) quantile greedy selection. For each method, we characterize both its theoretical computational complexity and practical applicability, specifying the problem scales where each variant proves to be most effective. These analyses provide system designers with clear guidelines for algorithm selection based on their specific latency requirements and workload characteristics.
@@
Numerical Experiments. In Section 7, we present comprehensive empirical evaluations using real-world data to validate our theoretical findings. To accurately capture the practical challenges of variable-length inputs, we construct a mixed dataset combining: (1) short conversational prompts from Zheng et al. (2023a), and (2) long document summarization tasks from Cohan et al. (2018). This experimental design reflects realistic deployment scenarios where LLM servers must handle diverse request types simultaneously. We evaluate four scheduling approaches: first-come-first-served, shortest-first, our LP-based Sorted-LP, and our main proposed algorithm Sorted-F. The results demonstrate that Sorted-F achieves significantly lower average latency, confirming both the theoretical robustness and practical effectiveness of our approach.
@@ 1.2 Other Related Works
In contrast, a few recent studies aim to develop theoretical foundations for LLM inference. The primary mathematical model considered in this paper builds on the framework introduced by Jaillet et al. (2025). Ao et al. (2025) examine inference scheduling under multiple objectives and propose algorithms with provable performance guarantees. Additionally, Chen et al. (2025, 2026) study the scheduling optimization with unknown decode lengths, and Li et al. (2025) analyze the stability conditions of LLM inference when operating on a single computational worker.
@@ 2 Model Review
In this section, we review the scheduling model introduced by Jaillet et al. (2025) for LLM inference under memory constraints. Before presenting the formal model, we provide background on two key features of modern LLM serving systems—prefill/decode mixing and continuous batching—that the model captures. We then discuss our extension, which relaxes a strong assumption on input sizes made in the original work.
Prefill/Decode (PD) mixing.
Modern LLM serving systems support prefill/decode (PD) mixing Kwon et al. (2023), whereby a single processing batch can simultaneously contain requests in different phases—some undergoing prefill while others are mid-decode. This mixing improves hardware utilization by combining diverse workloads within each batch.
Continuous Batching.
Traditional batching forms a fixed set of requests and processes them until all complete before admitting new requests. In contrast, continuous batching dynamically admits new requests into an ongoing batch as soon as memory capacity becomes available—typically when an in-progress request completes and releases its KV cache. This approach significantly improves throughput by avoiding idle capacity while waiting for long-running requests to finish.
Figure 1 illustrates the scheduling process. Circles represent prefill operations and squares represent decode operations, with colors distinguishing requests. Operations aligned vertically form a batch processed in parallel. The figure demonstrates PD mixing (e.g., prefill and decode operations co-occur in the same batch) and continuous batching (new requests enter as earlier ones complete and release memory).
Figure 1: Illustration of LLM inference scheduling with prefill/decode mixing and continuous batching. Circles denote prefill operations; squares denote decode tokens; colors distinguish requests. Operations in the same column constitute a batch processed simultaneously. The schedule exhibits PD mixing (batches contain both prefill and decode operations) and continuous batching (requests enter dynamically as memory becomes available).
2.1 Mathematical Model Review
@@
Jaillet et al. (2025) introduces a discrete-time system with a single computational worker. The worker has a KV cache capacity of $M>0$ tokens; in practice, $M$ depends on model size and hardware specifications (e.g., GPU memory) and is assumed known to the scheduler.
Request Arrivals and Adversarial Setting.
We consider $n$ prompts that arrive simultaneously at time $t=0$. Each prompt $i\in[n]$ is characterized by a pair $(s_{i},o_{i})$, where $s_{i}$ denotes the prefill size (number of input tokens) and $o_{i}$ denotes the number of output tokens to be generated. We adopt an adversarial framework: the adversary selects the number of requests $n$ and the parameters $(s_{i},o_{i})$ for each request, subject to a scaling constraint described below.
Scaling Assumption.
To state asymptotic guarantees as the memory budget $M$ grows, we impose a scaling constraint on the adversary. Define the peak per-request memory as
$L(M)\;:=\;\max_{i\in[n]}\{s_{i}+o_{i}\}.$
We require $L(M)=o(M)$, i.e., $L(M)/M\to 0$ as $M\to\infty$. This ensures that each individual request occupies a vanishing fraction of the memory budget, allowing concurrency to grow with hardware capacity.
Justification of the Scaling Assumption.
This assumption reflects practical deployment considerations. In real systems, operators first characterize their workload by estimating the rough distribution of prefill and decode lengths for expected requests. They then provision hardware with memory capacity $M$ substantially larger than typical request sizes to achieve high concurrency and throughput. Our scaling assumption $L(M)=o(M)$ captures precisely this regime where hardware is appropriately sized for the workload.
Output Length Information.
In the theoretical analysis, we assume that output lengths $o_{i}$ are known to the scheduler. This assumption isolates the fundamental scheduling challenge, which is already NP-hard (see Section 2.3). In practice, $o_{i}$ can be predicted, and several effective prediction methods exist (Zheng et al. 2023b, Qiu et al. 2024, Fu et al. 2024, Chen et al. 2024, Shahout et al. 2024). However, none of these prediction methods can guarantee accuracy. In Section 5, we introduce practical methodology that performs well even when prediction accuracy is very poor, rather than relying solely on point estimates.
Batch Processing and Memory Constraint.
Each request $i$ must be processed token by token during the decode phase. The memory required to generate the $j$-th output token ($j\in[o_{i}]$) is $s_{i}+j$, reflecting the growing KV cache. At each time step $t$, the system processes a batch $\mathcal{I}_{t}$ that may contain tokens from multiple requests at various stages of completion. Consistent with sequential decoding, at most one token from each request can appear in any batch. The model captures both PD mixing (batches combine requests in prefill and decode phases) and continuous batching (new requests may join at any time step when capacity permits). The memory constraint requires
@@
where $a_{i}$ is the index of the token from request $i$ in the batch.
Overloaded Regime and Justification of Offline Setting.
We restrict attention to overloaded instances where the total workload far exceeds single-batch capacity:
$\frac{n\,L(M)}{M}\;\longrightarrow\;\infty\qquad\text{as }M\to\infty.$
This restriction is essentially without loss of generality: when $n$ is small that all requests fit into a single batch, they can be processed simultaneously and scheduling becomes trivial. The algorithmic challenge arises precisely when requests must be sequenced across multiple time steps.
This overloaded regime is naturally consistent with the offline arrival model. In practice, LLM inference workloads exhibit tidal patterns: request volumes surge during peak hours and subside otherwise. During low-demand periods, requests are served immediately upon arrival, rendering scheduling unnecessary. During peak periods, however, arrivals exceed service capacity, creating a backlog of pending requests that must be prioritized. Our assumption that all $n$ prompts arrive simultaneously at $t=0$ models this backlogged state during peak demand, and our analysis targets this challenging regime.
2.2 Objective and Performance Metric
The objective is to minimize total end-to-end latency (TEL), defined as the sum of completion times. Since all requests arrive at $t=0$, the latency of request $i$ equals its completion time $c_{i}(\Lambda)$ under schedule $\Lambda$:
@@ 2.2 Objective and Performance Metric
Jaillet et al. (2025) formulate an integer linear program (ILP) to compute optimal schedules. However, solving the ILP is computationally prohibitive for real-time deployment, where decisions must be made within milliseconds. Our goal is therefore to design efficient approximation algorithms with provable performance guarantees.
Let Optimal denote an optimal schedule for request set $\mathcal{I}$. The performance of algorithm $\Lambda$ is measured by its competitive ratio:
@@
The competitive ratio is always at least 1, with smaller values indicating better worst-case performance.
2.3 Challenges with Heterogeneous Input Sizes
Jaillet et al. (2025) propose a memory-constrained shortest-first algorithm (MC-SF) that prioritizes requests by output length $o_{i}$. Before adding a request to the current batch, the algorithm verifies that memory constraints will remain satisfied at all future time steps.
Formally, let $\mathcal{R}^{(t)}$ denote requests not yet started by time $t$, and let $S^{(t)}$ denote requests currently in progress (with $p_{i}$ denoting the start time of request $i\in S^{(t)}$). For a candidate subset $U\subseteq\mathcal{R}^{(t)}$, define $t_{\max}(U):=\max_{i\in U}\{t+o_{i}\}$. The feasibility condition is:
$\sum_{i\in S^{(t)}}(s_{i}+t^{\prime}-p_{i})\cdot\mathbbm{1}_{\{o_{i}\geq t^{\prime}-p_{i}\}}+\sum_{i\in U}(s_{i}+t^{\prime}-t)\cdot\mathbbm{1}_{\{o_{i}\geq t^{\prime}-t\}}\leq M,\quad\forall t^{\prime}\in\{t,\ldots,t_{\max}(U)\}.$ (1)
Jaillet et al. (2025) establish strong theoretical and empirical performance of MC-SF under the assumption that $s_{i}=s$ for all $i\in[n]$. With uniform prefill sizes, shortest-first scheduling is natural: requests with smaller $o_{i}$ have lower peak memory usage ($s+o_{i}$) and shorter processing times, so prioritizing them improves both throughput and memory utilization.
However, uniform prefill sizes may not hold when systems handle diverse workloads mixing long and short prompts. Relaxing this assumption introduces fundamental difficulties. Consider requests $i$ and $j$ with $s_{i}<s_{j}$ but $o_{i}>o_{j}$: request $i$ has lower per-step memory but longer duration, while $j$ has higher memory but finishes quickly. The optimal priority depends on global interactions with other requests, making the trade-off nontrivial.
We show that MC-SF can perform arbitrarily poorly when input sizes vary.
@@ 2.3 Challenges with Heterogeneous Input Sizes
Suppose each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$, and $s_{i}+o_{i}\leq M$. Then MC-SF has unbounded competitive ratio: $\mathrm{CR}(\textup{{MC-SF}})\to\infty$ as $M\to\infty$.
The proof appears in Appendix 8, which also shows that prioritizing by total size $s_{i}+o_{i}$ yields an unbounded ratio. Moreover, the problem becomes computationally hard without uniform inputs. [^1] Whether minimizing TEL is NP-hard under uniform $s_{i}$ remains open.
@@
Under the model of Jaillet et al. (2025) with $s_{i}\in[1,M]$, $o_{i}\in[1,M]$, and $s_{i}+o_{i}\leq M$, minimizing total end-to-end latency is NP-hard.
We prove Theorem 2.2 via reduction from 3-Partition in Appendix 8.
@@ 3 Algorithm Description
where $\mathcal{X}$ is a set of requests, and $|\mathcal{X}|$ is the cardinality of the request set. Smaller values of $F(\mathcal{X})$ indicate higher scheduling priority. This metric fundamentally improves upon shortest-first methods by naturally balancing batch size and response lengths.
To illustrate its operational intuition, we begin with a concrete numerical example that demonstrates how Sorted-F outperforms shortest-first scheduling (MC-SF):
@@
Under MC-SF, which prioritizes shorter outputs, Type 1 would be scheduled first since $o_{1}=1<o_{2}=2$. After Type 1 completes at time $t=1$, Type 2 requests are processed in batches. With memory constraint $M=64$, each batch can accommodate $\lfloor 64/(1+2)\rfloor=21$ Type 2 requests simultaneously. The total end-to-end latency (TEL) is
@@
Building on this concrete demonstration, we now generalize the intuition through a symbolic example that shows how the F-metric recovers the optimal choice in a stylized setting: